OpenStudy (anonymous):

Solve for r 1400 = 1000 [(1+r/12) ^60] Will give medal. I'm stuck at 7/5 = (12+r)^60 ... don't know how to simplify that

OpenStudy (anonymous):

answer is ~ 0.0675 ....6.75% ...but I want to know how to simplify to get there.

OpenStudy (paxpolaris):

\[1400 = 1000\left( 1+\frac r {12} \right) ^{60}\]\[\frac75= \left( 1+\frac r {12} \right) ^{60}\] ...next take the log of both sides

OpenStudy (anonymous):

can you show those steps

OpenStudy (paxpolaris):

\[\log \left( \frac75 \right)=60\times \log \left( 1+\frac r {12} \right)\]

OpenStudy (paxpolaris):

\[{\log \left( 1.4 \right)\over 60}= \log \left( 1+\frac r {12} \right)\]

OpenStudy (anonymous):

okay, im following you. how do i get rid of the logs though?

OpenStudy (anonymous):

I need to get r by itself

OpenStudy (paxpolaris):

now reverse the log: \[10^{\log(1.4)/60}=1+r/12\]

OpenStudy (anonymous):

that doesn't get the right answer

OpenStudy (anonymous):

oh wait it does!

OpenStudy (anonymous):

that becomes 10 to the power of 0.0024....and then multiply that by 12

OpenStudy (anonymous):

which there was an easier way of doing it then reversing the log

OpenStudy (anonymous):

wish*